# Classical Simulations of Low Magic Quantum Dynamics

### v1

@@ I Introduction
The amount of non-Clifford resources required to prepare a state is known as non-stabilizerness, commonly referred to as “magic”. Magic is zero for stabilizer states and increases monotonically under the application of non-Clifford operations. In recent years, a range of near-stabilizer simulation methods have been developed to efficiently simulate quantum circuits with limited non-Clifford resources. These include Monte Carlo techniques based on stabilizer decompositions [11, 16], as well as tensor network methods such as MPS simulations [17]. Additionally, hybrid approaches such as stabilizer tensor networks, which combine stabilizer formalism with tensor network contraction and sampling, have been developed  [18]. Collectively, these techniques enable approximate classical simulation of quantum circuits that are not fully Clifford but remain structured enough to avoid worst-case exponential scaling of simulation complexity with the number of non-Clifford gates.
@@
The remainder of this paper is organized as follows. In Section II, we introduce our algorithm for updating near-stabilizer states under the action of $T$-gates. In Section IV, we define the measure of magic, stabilizer nullity, which we study in this work. In Sections III, we introduce the models of monitored circuits, including entangling gates and protocols for the measurements. In Section IV.1, we test the accuracy of our algorithm, which simulates the full density matrix, by varying a cutoff parameter that controls the accuracy of the algorithm. We also identify the magic transition in these circuits. In Section II.2, we examine the space complexity of these circuits, namely how the amount of memory to store the state scales with the density of $T$-gates. In Section B, we study the entanglement transitions in these models, focusing on the purification transition. In Appendix A, we provide more details on the algorithms used to study these circuits. These include the evolution of the density matrix of the generic state $|\psi\rangle$, $\rho=|\psi\rangle\langle\psi|$, under measurements and Clifford unitaries, computing the Rényi entanglement entropies, and Bell sampling. In Appendix B, we examine entanglement transitions in more detail. Namely, we locate cluster and all-to-all entanglement transitions. We also provide more details on locating the purification transition. In Appendix C, we include details on the procedure used to locate the magic transition.
@@ II.1 Stabilizer Formalism
The density matrix of a general mixed stabilizer state on $L$ qubits can be written as a sum over the elements of its stabilizer group $\mathcal{S}$:
@@
Figure 1: (a) Circuit diagram for the evolution of initial state $\rho_{o}=\sum_{\ell}\lambda_{\ell}\sigma_{\ell}\rho_{S}$ with measurements, $T$-gates and Clifford unitaries (b) Schematic of the forking of the logicals $\sigma_{\ell}$ for the initial state under the circuit evolution in (a). The LRSD is represented diagrammatically, where each node represents $\sigma_{\ell}$ or $\lambda_{\ell}.$ Vertical edges connect terms from the initial state $\rho_{o}$ to the final state after the T-gate $T\rho_{o}T^{\dagger}$, or measurement of Pauli operator $P$, $\frac{I+P}{2}\rho_{o}\frac{I+P}{2}$. Each $\sigma_{\ell}$ is mapped to a new set of $\sigma_{\ell}$ attached to it by vertical lines. An X, or box with T, on a branch denotes measurement, or $T$-gate, respectively, in the entire circuit. (c) Equation for the LRSD and the effect of $T$-gates and measurements. (d) Each injection of a $T$-gate can fork each logical onto itself (Case I), all logicals into three new logicals (Cases II, IV), or certain logicals to three new logicals (Case III). The Pauli operator $\overline{P}$ is defined in Eq. (9). The Pauli $\overline{P}_{(X/Y)}$ is defined by $\overline{P}_{X/Y}=\overline{P}_{1}\otimes\cdots\otimes(X/Y)\otimes\cdots\otimes\overline{P}_{L},$ where $X/Y$ and the $T$-gate is on the $i^{\mathrm{th}}$ qubit and likewise $\sigma_{\ell}^{(X/Y)}=\sigma_{\ell_{1}}\otimes\cdots\otimes(X/Y)\otimes\cdots\otimes\sigma_{\ell_{L}}.$ (e) Each measurement can fork each logical onto itself (Case I), eliminate all anticommuting logicals (Cases II), or update anticommuting logicals with the Pauli operator $\overline{P}$. Algorithm 1 in Appendix A provides details on measurement updates.
@@ II.2 Non-Clifford Evolution
The low-rank stabilizer decomposition (LRSD) can represent the density matrix of an arbitrary pure quantum state $|\psi\rangle$ as
@@
Here, $\rho_{S}$ is a stabilizer state which may be mixed or pure. The coefficients $\lambda_{l}$ are real numbers and the logical operators $\sigma_{l}$ are defined in Eq (4). For states with a low stabilizer rank $\chi$, the number of logical operators needed in the LRSD remains small, allowing for an efficient classical representation.
This approach isolates the non-stabilizer content of the state into a compact, low-dimensional structure, the logical operator set, while leveraging the efficient classical simulability of stabilizer states. For states with low stabilizer rank, the decomposition remains compact, making the LRSD an efficient and scalable representation for near-Clifford states.
The logical operators, or logicals ($\sigma_{\ell}$), are indexed by $\ell$ and defined as
$\sigma_{\ell}\;=\;\bigotimes_{j=1}^{L}Q_{\ell j},\qquad Q_{\ell j}\in\{I,X,Y,Z\},$ (4)
Let us define the logical set $\mathcal{L}_{\mathrm{LRSD}}=\{\sigma_{l}\}$ as a collection of Pauli operators that form a vector space whose field is real numbers and operation is matrix addition:
$\mathrm{span}(\mathcal{L}_{\mathrm{LRSD}})=\left\{\sum_{i}\lambda_{i}\sigma_{i}\;\middle|\;\sigma_{i}\in\mathcal{L}_{\mathrm{LRSD}},\;\lambda_{i}\in\mathbb{R}\right\}.$ (5)
Any Hermitian operator $\mathcal{O}$ within this space can be expressed as a real linear combination of elements in $\mathcal{L}_{\mathrm{LRSD}}$. We use this structure to specify a pure quantum state $|\psi\rangle$ in terms of a fixed mixed stabilizer density matrix, $\rho_{S}$, along with a set of logical Pauli operators $\mathcal{L}_{\mathrm{LRSD}}$ and real weights $\{\lambda_{l}\}$. While this decomposition is not unique, it provides a practical and scalable method for simulating low-rank stabilizer states. We now present how to update this representation of the state under the action of $T$-gates. For a $T$-gate on the $i^{th}$ qubit, $T_{i}$, the state evolves as:
@@
    $T_{i}\sigma_{q}T_{i}^{\dagger}=(\alpha^{2}+\beta^{2})\sigma_{q}+2i\alpha\beta\epsilon_{\text{qr3}}\sigma_{r}.$ (11)
    Here, $\epsilon_{\text{qr3}}$ is the Levi-Civita symbol and $\sigma_{q}$ is a single-qubit Pauli matrix, where each index runs from $1$ to $3$.
@@
    Each term in the tensor product of $\sigma_{l}$ is updated using Eq. 11.
@@
The treatment of Clifford unitaries and Pauli measurements appears in Appendix A. We can now provide some intuition for how the complexity of the classical simulation evolves. Each $T$-gate application can increase the number of logicals $|\mathcal{L}_{\mathrm{LRSD}}|$. Specifically, in Cases II or III, the number of logicals can triple. Conversely, each measurement can decrease the number of logicals by a factor of 1 if $P,\mathcal{L}_{\mathrm{LRSD}}]=0$ or $P,\mathcal{L}_{\mathrm{LRSD}}]\neq 0$ and $P,\mathcal{S}]\neq 0$ or three if $P,\mathcal{L}_{\mathrm{LRSD}}]\neq 0$ and $P,\mathcal{S}]=0.$ The details of the procedure are given in Algorithm 3.
@@
We now define a measure of magic called the stabilizer nullity [25]. If $|\psi\rangle$ is an $L$ qubit state, the stabilizer group $\mathcal{S}$, of $|\psi\rangle$ is
where $\mathcal{P}_{L}$ is the Pauli group on $L$ qubits and $\mathcal{S}$ is a subgroup of $\mathcal{P}$. The stabilizer nullity of $|\psi\rangle$, $\mathcal{M}(|\psi\rangle)$, is
@@
where $|\absolutevalue{\mathcal{S}(|\psi\rangle)}$ is the size of the stabilizers. For a stabilizer state, $\mathcal{M}=0.$ For non-stabilizer states, $\mathcal{M}>0.$ The stabilizer nullity measures the number of non-stabilizer degrees of freedom in the state. It serves as an efficiently computable lower bound on more general measures of magic, such as stabilizer rank or robustness of magic [12, 26, 27, 28].
@@ III Circuit Model
Figure 2: (a) Circuit diagram for the single-pair all-to-all model. Each time step consists of a CZ gate applied between two random sites, followed by a measurement on a randomly chosen site in either the $X$ or $Z$ basis. Each dashed line represents the end of a single time step in the model. (b) Forking of the initial stabilizer wavefunction $|\psi\rangle=|+\rangle^{\otimes L}$ under circuit evolution with $T$-gates, measurements, and Clifford unitaries. Each injection of $T$-gate takes a stabilizer state $|\psi\rangle$ onto a superposition of at most two stabilizer states $|\psi^{\prime}\rangle=c_{+}|\psi_{+}\rangle+c_{-}|\psi_{-}\rangle.$ Each branch represents a different stabilizer state. Measurements can collapse different branches onto a single branch. Each quantum trajectory is represented by a different color. forked
In all-to-all random circuit models, there is no constraint of spatial locality. Part of the motivation of studying these models is that a graphical description of the MIPT in these circuits leads to a precise connection between the MIPT and mean-field percolation [29]. This is analogous to the connection between two-dimensional classical percolation and the measurement-driven entanglement transition in one spatial dimension [6]. Furthermore, all-to-all models are directly relevant to various experimental platforms. For example, proposals for scaling up quantum computers often involve quantum network architectures, where locally interacting nodes are connected via non-local photonic interconnects, leading to effectively all-to-all connectivity at large scales [30]. At the hardware level, trapped ion quantum computers already realize all-to-all gate connectivity among physical qubits at intermediate scales [31, 32]. These experimental developments further underscore the importance of understanding the fundamental features of MIPTs in nonlocal circuit architectures.
We now introduce the circuit protocols considered in this paper. We focus on an all-to-all geometry where the two-site unitary can be applied to any pair of qubits, which is an effective 0D system. Specifically, we consider single-pair all-to-all models, where in a single time step, at most one gate couples a pair of qubits, as shown in Fig. 2. Each time step consists of three probabilistic stages: Clifford unitaries, non-Clifford unitaries, and measurements.
@@ IV.1 Magic Complexity
To test the accuracy of the algorithm, with a truncation error threshold $\epsilon$, we evolve the LRSD under the circuit, truncated at each timestep to keep only the coefficients $\lambda_{l}$ such that $|\lambda_{l}|>\epsilon$. If there are no remaining logicals with $|\lambda_{l}|>\epsilon$, the trajectory is invalid and not considered. We pick $\epsilon$ sufficiently small so that this issue does not arise. In Fig. 3, we plot the dependence of the total number of logicals, $\overline{|\{\lambda_{l}\}|}$ on $\epsilon$ for varying $p_{m,Z}$. The $\overline{\cdots}$ denotes an average over circuit trajectories and $|\{\cdots\}|$ is the set size. We fix the location of the unitary gates and measurement locations to ensure $\overline{\left|\{\lambda_{l}\}\right|}(\epsilon)$ depends only on the truncation error and not the randomness of the circuit-to-circuit fluctuations. As shown in Fig. 3, as the cutoff is varied from $\epsilon=0$ to $\epsilon\sim 10^{-2}$, the number of logicals is unchanged, which demonstrates that for this given circuit, this is the required number of logicals. In contrast, further increasing the cutoff to $\epsilon\sim\mathcal{O}(1)$, $|\{\lambda_{l}\}|$ is reduced by roughly two orders of magnitude, $\mathcal{O}(10^{2})$, in the effective Hilbert space dimension, shown in 3(a). This greatly reduces the computational cost of Bell sampling at the price of a less accurate wavefunction. We also clearly see the role of $p_{m,Z}$ measurements, which reduces the number of required logicals (a) and we find that (b) the stabilizer nullity grows by no more than five percent across the entire measurement rate window $p_{m,Z}\in\{0.6,0.7,0.8,0.9\}$ as compared to the zero truncation curves (dashed lines). These observations demonstrate that non-trivial observables remain robust to moderate truncation of the LRSD. This allows us to introduce $\epsilon$ as an effective cutoff scheme to reduce simulation time while preserving quantitative accuracy.
@@
Following the Algorithm in Sec II.2 for applying $T$-gates, each $T$-gate will at most triple $\overline{\left|\{\lambda_{l}\}\right|}.$ This means, in the worst case scenario, $\overline{\left|\{\lambda_{l}\}\right|}=3^{|T|}$, where $|T|$ is the number of $T$-gates in the circuit. In this section, we consider a model where a single layer of $T$-gates is applied at the final timestep of the Clifford circuit. That is, we sweep from $i=1$ to $i=L$ once on a Clifford state. This means $\overline{\left|\{\lambda_{l}\}\right|}$ could, at worst, grow to $3^{L}$. The origin of this $3^{L}$ scaling comes from the fact that each of the $L$ $T$-gates can, in the worst case, triple the number of terms in the LRSD. The rapid growth in $\overline{\left|\{\lambda_{l}\}\right|}$ imposes a challenge in the classical simulation. Therefore, only the state within the area law magic phase can be efficiently simulated using LRSD. In the area law magic phase, we find that the number of logicals, when averaged over circuits, scales with system size as a power law, as shown in Figure 4.
@@
Figure 3: (a) $\overline{\left|\{\lambda_{l}\}\right|}$ as a function of truncation error $\epsilon$ for $\epsilon$ in $10^{-4},0.5]$. The $\overline{\phantom{X}}$ denotes averaging over circuits. $|\{\lambda_{l}\}|$ represents the total number of logicals $\lambda_{l}$, where $\lambda_{l}$ is defined in Eq. 9. Data is shown for the single-pair all-to-all model with Z-basis measurements, where Bell Sampling is performed on $|\psi\rangle\otimes|\psi\rangle$ for system size $L=50$, probability of $T$-gates $p_{T}=1/L$, and measurement probabilities $p_{m,Z}\in\{0.6,0.7,0.8,0.9\}$, with a darker color denoting a larger measurement probability. (b) Stabilizer nullity $\overline{\mathcal{M}}$, as defined in Eq. (64), as a function of the cutoff $\epsilon$ for several measurement probabilities $p_{m,Z}$. The dashed line denotes data obtained at zero truncation error. Data was obtained by averaging over $\mathcal{O}(10^{3})$ circuit realizations. For each set of $\epsilon$ at a fixed $p_{m,Z}$, the locations of unitary gates and measurements are fixed.
@@
Figure 4: (a) Scaling of the circuit-averaged number of entries $\overline{\mathcal{N}}$ with $L$ (see Eq. (21)) is the total number of entries required to specify the state in our algorithm. $|\{\lambda_{l}\}|$ represents the total number of logicals $\lambda_{l}$, where $\lambda_{l}$ is defined in Eq. 9. Data is shown for the single-pair all-to-all model with X-basis measurements. For the statevector simulation, $\mathcal{\overline{N}}=2^{L}$, as shown by the black dashed line. Data is shown for measurement probabilities $p_{m,X}\in\{0.6,0.9\}$ (b) Data is shown for cutoff values $\epsilon\in\{1\times 10^{-15},1\times 10^{-10},1\times 10^{-6},1\times 10^{-5},5\times 10^{-4}\}$, with a darker color denoting a larger cutoff. (c) Data is shown for $T$-gate probabilities $p_{T}\in\{1/L,0.75/L,0.5/L,0.25/L\}$. Data was obtained by averaging over $\mathcal{O}(10^{3})$ circuit realizations.

### v2

@@ I Introduction
The amount of non-Clifford resources required to prepare a state is known as non-stabilizerness, commonly referred to as “magic”. Magic is zero for stabilizer states and increases monotonically under the application of non-Clifford operations. In recent years, a range of near-stabilizer simulation methods have been developed to simulate quantum circuits with limited non-Clifford resources efficiently. These include Monte Carlo techniques based on stabilizer decompositions [11, 16, 17], mapping to classical statistical mechanics models [18] as well as tensor network methods such as MPS simulations [19] and Clifford-augmented Matrix Product States (CAMPS) [20, 21]. Additionally, hybrid approaches such as stabilizer tensor networks, which combine stabilizer formalism with tensor network contraction and sampling, have been developed  [22]. Collectively, these techniques enable approximate classical simulation of quantum circuits that are not entirely Clifford but remain structured enough to avoid worst-case exponential scaling of simulation complexity with the number of non-Clifford gates.
@@
The remainder of this paper is organized as follows. In Section II, we introduce our algorithm for updating near-stabilizer states under the action of $T$-gates. We also define the measure of magic, stabilizer nullity, which we study in this work. In Section III, we introduce the models of monitored circuits, including entangling gates and protocols for the measurements. In Section IV.1, we test the accuracy of our algorithm, which simulates the full density matrix, by varying a cutoff parameter that controls the accuracy of the algorithm. In Section IV.2, we examine the space complexity of these circuits, namely how the amount of memory to store the state scales with the density of $T$-gates. In Section V, we study the MIPTs in these models, focusing on the magic and purification transitions. In Appendix A, we provide more details on the algorithms used to study these circuits. These include the evolution of the density matrix of the generic state $|\psi\rangle$, $\rho=|\psi\rangle\langle\psi|$, under measurements and Clifford unitaries, computing the Rényi entanglement entropies, and Bell sampling. In Appendix B, we examine entanglement transitions in more detail. Namely, we locate cluster and all-to-all entanglement transitions. We also provide more details on locating the purification transition. In Appendix C, we include details on the procedure used to locate the magic transition.
@@ II.1 Stabilizer Formalism
The density matrix of a general mixed stabilizer state on $L$ qubits can be written as a sum over the elements of its stabilizer group $\mathcal{S}$ [10]:
@@
Figure 1: (a) Circuit diagram for the evolution of the initial state that corresponds to the initial density matrix $\rho_{o}=\sum_{\ell}\lambda_{\ell}\sigma_{\ell}\rho_{S}$ with measurements, $T$-gates and Clifford unitaries (b) Schematic of the forking of the logicals $\sigma_{\ell}$ for the initial state under the circuit evolution in (a). The LRSD is represented diagrammatically, where each node represents $\sigma_{\ell}$ or $\lambda_{\ell}.$ Vertical edges connect terms from the initial state $\rho_{o}$ to the final state after the T-gate $T\rho_{o}T^{\dagger}$, or measurement of Pauli operator $P$, $\frac{I+P}{2}\rho_{o}\frac{I+P}{2}$. Each $\sigma_{\ell}$ is mapped to a new set of $\sigma_{\ell}$ attached to it by vertical lines. An X, or box with T, on a branch denotes measurement, or $T$-gate, respectively, in the entire circuit. (c) Equation for the LRSD and the effect of $T$-gates and measurements. (d) Each injection of a $T$-gate can fork each logical onto itself (Case I), all logicals into three new logicals (Cases II, IV), or certain logicals to two new logicals (Case III). The Pauli operator $\overline{P}$ is defined in Eq. (8). The Pauli $\overline{P}_{(X/Y)}$ is defined by $\overline{P}_{X/Y}=\overline{P}_{1}\otimes\cdots\otimes(X/Y)\otimes\cdots\otimes\overline{P}_{L},$ where $X/Y$ and the $T$-gate is on the $i^{\mathrm{th}}$ qubit and likewise $\sigma_{\ell}^{(X/Y)}=\sigma_{\ell_{1}}\otimes\cdots\otimes(X/Y)\otimes\cdots\otimes\sigma_{\ell_{L}}.$ (e) Each measurement can fork each logical onto itself (Case I), eliminate all anticommuting logicals (Cases II), or update anticommuting logicals with the Pauli operator $\overline{P}$. Algorithm 1 in Appendix A provides details on measurement updates.
@@ II.2 Non-Clifford Evolution
Given an arbitrary pure quantum state $|\psi\rangle$, the stabilizer group of $|\psi\rangle$ is the set of Pauli operators that have eigenvalue one on that state
where $\mathcal{P}_{L}$ is the Pauli group on $L$ qubits and $\mathcal{S}$ is a subgroup of $\mathcal{P}_{L}$. We define the low-rank stabilizer decomposition (LRSD) of $|\psi\rangle$ as the representation
@@
Here, $\rho_{S}$ is a (possibly mixed) stabilizer state that is defined solely in terms of $S$ through Eq. (1). The coefficients $\lambda_{l}=\langle\psi|\sigma_{l}|\psi\rangle$ are real numbers and the “logical” operators $\sigma_{l}$ are elements of the Pauli group that commute with $S$. The set of all logical operators modulo phases $\mathcal{L}_{\mathrm{LRSD}}$ has size $4^{k}$ for a stabilizer group of size $2^{L-k}$ [10]. If $\rho_{S}$ is pure, then $|\psi\rangle$ contains no magic, and Eq. (4) consists of a single term: specifically, $\lambda_{l}=1$ and $\sigma_{l}=I$. In this case, $|\psi\rangle$ is a stabilizer state. More generally, for a broad class of near-stabilizer states with small $k$, the number of logical operators needed in the LRSD remains small, allowing for an efficient classical representation. This approach isolates the non-stabilizer content of the state into a compact, low-dimensional structure, the logical operator set, while leveraging the efficient classical simulability of stabilizer states.
We now present how to update the LRSD representation of the state under the action of $T$-gates. For a $T$-gate on the $i^{th}$ qubit, $T_{i}$, the state evolves as:
@@
    $T_{i}\sigma_{q}T_{i}^{\dagger}=(\alpha^{2}+\beta^{2})\sigma_{q}+2i\alpha\beta\epsilon_{qr3}\sigma_{r}.$ (10)
    Here, $\epsilon_{qr3}$ is the Levi-Civita symbol and $\sigma_{q}$ is a single-qubit Pauli matrix, where each index runs from $1$ to $3$.
@@
    Each term in the tensor product of $\sigma_{l}$ is updated using Eq. (10).
@@
The treatment of Clifford unitaries and Pauli measurements appears in Appendix A. We can now provide some intuition for how the complexity of the classical simulation evolves. Each $T$-gate application can increase the number of logicals $|\mathcal{L}_{\mathrm{LRSD}}|$. Specifically, in Cases II and IV, the number of logicals can triple, while in Case III, they can double. Conversely, each measurement can reduce the number of logicals, depending on the commutation relationships. If $P,\mathcal{L}_{\mathrm{LRSD}}]=0$, or if $P,\mathcal{L}_{\mathrm{LRSD}}]\neq 0$ and $P,\mathcal{S}]\neq 0$, the number of logicals remains the same. However, if $P,\mathcal{L}_{\mathrm{LRSD}}]\neq 0$ and $P,\mathcal{S}]=0$, the number of logicals can decrease by up to a factor of three. The detailed procedure is provided in Algorithm 3.
@@
We now define a measure of magic called the stabilizer nullity [29]. If $|\psi\rangle$ is an $L$ qubit state, the stabilizer group $\mathcal{S}$, of $|\psi\rangle$ is defined in Eq. (3) The stabilizer nullity of $|\psi\rangle$, $\mathcal{M}(|\psi\rangle)$, is
@@
where $|\absolutevalue{\mathcal{S}(|\psi\rangle)}$ is the size of the stabilizers. For a stabilizer state, $\mathcal{M}=0.$ For non-stabilizer states, $\mathcal{M}>0.$ The stabilizer nullity measures the number of non-stabilizer degrees of freedom in the state. It serves as an efficiently computable upper bound on more general measures of magic, such as stabilizer rank or robustness of magic [12, 30, 31, 32].
@@ III Circuit Model
Figure 2: (a) Circuit diagram for the single-pair all-to-all model. Each time step consists of a CZ gate applied between two random sites, followed by a measurement on a randomly chosen site in either the $X$ or $Z$ basis. Each dashed line represents the end of a single time step in the model. (b) Forking of the initial stabilizer wavefunction $|\psi\rangle=|+\rangle^{\otimes L}$ under circuit evolution with $T$-gates, measurements, and Clifford unitaries. Each injection of $T$-gate takes a stabilizer state $|\psi\rangle$ onto a superposition of at most two stabilizer states $|\psi^{\prime}\rangle=c_{+}|\psi_{+}\rangle+c_{-}|\psi_{-}\rangle.$ Each branch represents a different stabilizer state. Measurements can collapse different branches onto a single branch. Each quantum trajectory is represented by a different color.
In all-to-all random circuit models, there is no constraint of spatial locality. Part of the motivation of studying these models is that a graphical description of the MIPT in these circuits leads to a precise connection between the MIPT and mean-field percolation [33]. This is analogous to the connection between two-dimensional classical percolation and the measurement-driven entanglement transition in one spatial dimension [6]. Furthermore, all-to-all models are directly relevant to various experimental platforms. For example, proposals for scaling up quantum computers often involve quantum network architectures, where locally interacting nodes are connected via non-local photonic interconnects, leading to effectively all-to-all connectivity at large scales [34]. At the hardware level, trapped ion quantum computers already realize all-to-all gate connectivity among physical qubits at intermediate scales [35, 36]. These experimental developments further underscore the importance of understanding the fundamental features of MIPTs in nonlocal circuit architectures. We now introduce the circuit protocols considered in this paper. We focus on an all-to-all geometry where the two-site unitary can be applied to any pair of qubits, which is an effective 0D system. Specifically, we consider single-pair all-to-all models, where in a single time step, at most one gate couples a pair of qubits, as shown in Fig. 2. Each time step consists of three probabilistic stages: Clifford unitaries, non-Clifford unitaries, and measurements.
@@ IV.1 Magic Complexity
To test the accuracy of the algorithm, with a truncation error threshold $\epsilon$, we evolve the LRSD under the circuit, truncated at each timestep to keep only the coefficients $\lambda_{l}$ such that $|\lambda_{l}|>\epsilon$. If there are no remaining logicals with $|\lambda_{l}|>\epsilon$, the trajectory is invalid and not considered. We pick $\epsilon$ sufficiently small so that this issue does not arise.
@@
In Fig. 3, we plot the dependence of the total number of logicals, $\overline{|\{\lambda_{l}\}|}$ on $\epsilon$ for varying $p_{m,Z}$. The $\overline{\cdots}$ denotes an average over circuit trajectories and $|\{\cdots\}|$ is the set size. We fix the location of the unitary gates and measurement locations to ensure $\overline{\left|\{\lambda_{l}\}\right|}(\epsilon)$ depends only on the truncation error and not the randomness of the circuit-to-circuit fluctuations. As shown in Fig. 3, as the cutoff is varied from $\epsilon=0$ to $\epsilon\approx 10^{-2}$, the number of logicals is unchanged, which demonstrates that for this given circuit, this is the required number of logicals. In contrast, further increasing the cutoff to $\epsilon\sim\mathcal{O}(1)$, $|\{\lambda_{l}\}|$ is reduced by roughly two orders of magnitude, $\mathcal{O}(10^{2})$, in the effective Hilbert space dimension, shown in Fig. 3(a). This greatly reduces the computational cost of Bell sampling at the price of a less accurate wavefunction. We also clearly see the role of $p_{m,Z}$ measurements, which reduces the number of required logicals (a) and we find that (b) the stabilizer nullity grows by no more than five percent across the entire measurement rate window $p_{m,Z}\in\{0.6,0.7,0.8,0.9\}$ as compared to the zero truncation curves (dashed lines). These observations demonstrate that non-trivial observables remain robust to moderate truncation of the LRSD. This allows us to introduce $\epsilon$ as an effective cutoff scheme to reduce simulation time while preserving quantitative accuracy.
@@
Following the Algorithm in Sec II.2 for applying $T$-gates, each $T$-gate will at most triple $\overline{\left|\{\lambda_{l}\}\right|}.$ This means, in the worst case scenario, $\overline{\left|\{\lambda_{l}\}\right|}=3^{|T|}$, where $|T|$ is the number of $T$-gates in the circuit. In this section, we consider a model where a single layer of $T$-gates is applied at the final timestep of the Clifford circuit. That is, we sweep from $i=1$ to $i=L$ once on a Clifford state. This means $\overline{\left|\{\lambda_{l}\}\right|}$ could, at worst, grow to $3^{L}$. The origin of this $3^{L}$ scaling comes from the fact that each of the $L$ $T$-gates can, in the worst case, triple the number of terms in the LRSD. The rapid growth in $\overline{\left|\{\lambda_{l}\}\right|}$ imposes a challenge in the classical simulation. Therefore, only the state within the area law magic phase can be efficiently simulated using LRSD. In the area law magic phase, we find that the number of logicals, when averaged over circuits, scales with system size as a power law, as shown in Fig. 4.
@@
Figure 3: (a) $\overline{\left|\{\lambda_{l}\}\right|}$ as a function of truncation error $\epsilon$ for $\epsilon$ in $10^{-4},0.5]$. The $\overline{\cdots}$ denotes averaging over circuit trajectories. $|\{\lambda_{l}\}|$ represents the total number of logicals $\lambda_{l}$, where $\lambda_{l}$ is defined in Eq. (4). Data is shown for the single-pair all-to-all model with Z-basis measurements, where Bell Sampling is performed on $|\psi\rangle\otimes|\psi\rangle$ for system size $L=50$, probability of $T$-gates $p_{T}=1/L$, and measurement probabilities $p_{m,Z}\in\{0.6,0.7,0.8,0.9\}$, with a darker color denoting a larger measurement probability. (b) Stabilizer nullity $\overline{\mathcal{M}}$, as defined in Eq. (13), as a function of the cutoff $\epsilon$ for several measurement probabilities $p_{m,Z}$. The dashed line denotes data obtained at zero truncation error. Data was obtained by averaging over $\mathcal{O}(10^{3})$ circuit realizations. For each set of $\epsilon$ at a fixed $p_{m,Z}$, the locations of unitary gates and measurements are fixed.
@@
Figure 4: (a) Scaling of the circuit-averaged number of entries $\overline{\mathcal{N}}$ with $L$, see Eq. (19), which is the total number of entries required to specify the state in our algorithm. $|\{\lambda_{l}\}|$ represents the total number of logicals $\lambda_{l}$, where $\lambda_{l}$ is defined in Eq. (4). Data is shown for the single-pair all-to-all model with X-basis measurements. For the statevector simulation, $\mathcal{\overline{N}}=2^{L}$, as shown by the black dashed line. Data is shown for measurement probabilities $p_{m,X}\in\{0.6,0.9\}$ (b) Data is shown for cutoff values $\epsilon\in\{1\times 10^{-15},1\times 10^{-10},1\times 10^{-6},1\times 10^{-5},5\times 10^{-4}\}$, with a darker color denoting a larger cutoff. (c) Data is shown for $T$-gate probabilities $p_{T}\in\{1/L,0.75/L,0.5/L,0.25/L\}$. Data was obtained by averaging over $\mathcal{O}(10^{3})$ circuit realizations.